### v1

@@ I Introduction
The amplitude of the nuclear M1 transition is suppressed by five orders of magnitude compared to typical atomic E1 transitions. However, it is possible to enhance this transition using the electronic bridge (EB) process, where electronic transitions induce nuclear transitions via hyperfine interaction (hfi). The EB effect also increases the probability of decay of the nuclear excited state. Corresponding calculations have been performed for Th II, Th III, and Th IV ions [14, 15, 16, 17, 18, 19, 20, 21].
Previous EB calculations for the decay [15] and excitation [16] of the excited nuclear state in the Th II ion could not provide accurate results due to large errors in the measured value of the nuclear excitation energy and the absence of measurements of the positions of Th II electron energy levels near the nuclear excitation energy. According to theoretical expectations [22], the spacing between electron energy levels in this region is very small, making resonance with the nuclear excitation possible.
@@
II Electronic bridge for nuclear excitation and decay in Th II
We follow the method used in Refs. [16, 21] for the Th II anf Th III ions. The decay of any atomic state with energy $E_{n}>\omega_{N}$ may include nuclear excitation. Since the nuclear excitation energy $\omega_{N}=67393$ cm-1 lies outside the optical region, we consider a two-step excitation of the electronic states(see also [16]). In the first step, the atom is excited from the ground state (GS) to one of the intermediate states $t$ used in [16]. In the second step, the ion is further excited by a second laser with frequency which is chosen to satisfy the energy conservation condition for simultaneous nuclear excitation and excitation of the final electronic state, $\omega_{1}+\omega_{2}=\omega_{N}+E_{s}$ ($\omega_{1}=E_{t}$). Nuclear excitation is due to the hyperfine interaction between atomic electrons and the nucleus. Corresponding diagram is presented on Fig. 1. Its calculation involves a summation over all intermediate states $n$. However, in this analysis, we focus on the dominant contributions, which come from specific intermediate states $n$ and final states $s$ providing a very small energy denominator in Eq. (3). Note that the final state $s$ is not the ground state but a low-lying excited state satisfying selection rules for the hfi-induced E1 transition from the state $t$ (see diagram on Fig. 1). We do not go to highly excited states $s$ to keep laser frequencies within optical region. For each final electron state $s$, we look for an intermediate state $n$ (using energy level data from Ref. [23]) which corresponds to the smallest energy denominator $\Delta$ in (3), $\Delta=E_{n}-E_{s}-\omega_{N}$, and use this state for estimation of the EB amplitude.
@@ III Numerical calculations of the electronic bridge contribution to the nuclear excitation
Table 1: Versions of the EB process with resonant nuclear excitation. The notations $t$, $n$, and $s$ refer to the diagram on Fig. 1. It is assumed that $E_{t}$ = 36390.53 cm-1, $\omega_{2}=\omega_{N}+E_{s}-E_{t}$ is the frequency of second laser excitation. The energy denominator in Eq. (3) is given by $\Delta=E_{n}-E_{s}-\omega_{N}$. The third column lists all possible values of $J_{n}$ from the experimental data in Ref. [23], the last two columns present the calculated values of the EB enhancement factors $\beta$ for specific $J_{n}$. Here, $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$, and $\langle x\rangle$ represents the average value of $x$ over theoretical states that are candidates to represent the actual states $t$ and $n$. No averaging is performed over $s$ since it is a low-lying state that is easily identified in the calculations. The median value of $x$ is denoted by $x_{m}$. Numbers in square brackets indicate powers of ten.
$N$ $E_{n}$ [^1] Ref. [23], Supplemental Material. $J_{n}$ $E_{s}$222 NIST database, [24]. $J_{s}$ $\omega_{2}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
@@
1 73637.54 3/2 or 5/2 6244.29 1/2 37247.10 -0.09 5.5[6]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole hfi. 1.6[5]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole hfi.
2 73637.54 3/2 or 5/2 6244.29 1/2 37247.10 -0.09 1.2[4]444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole hfi contributes. 660444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole hfi contributes.
3 74396.40 3/2 or 5/2 7001.42 3/2 38004.23 1.64 2.5[4] 800
4 76788.94 1/2 or 3/2 9400.96 5/2 40403.77 -5.36 5.6[4]555 if $J_{n}$=3/2. 700555 if $J_{n}$=3/2.
5 76895.40 3/2 9400.96 5/2 40403.77 101.1 3.0[3] [^6] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1. 4666Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1.
6 76445.96 1/2 or 3/2 9061.10 5/2 40063.91 -8.48 7.1[3]555 if $J_{n}$=3/2. 24555 if $J_{n}$=3/2.
7 76508.84 3/2 - 7/2 9061.10 5/2 40063.91 54.4 1.1[3] [^7] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1. 4777Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1.
8 68921.30 3/2 or 5/2 1521.90 5/2 32524.71 6.06 2.6[3] 28
9 75840.96 3/2 or 5/2 8460.35 5/2 39463.16 -12.73 5.9[3] 94
@@
In the first step, we identified the final electronic state $s$ and the intermediate electronic level $n$ that provide very small energy denominators $\Delta=E_{n}-E_{s}-\omega_{N}$ in Eq. (3). The smallest value of $\Delta$ is $-0.09$ cm-1. We also considered five additional cases with larger $\Delta$, where enhanced electronic bridge (EB) contributions may still be expected.
@@
For each final state $s$, we calculated 217 possible values of $x$, corresponding to different assignments of the experimental levels $t$ and $n$. The distribution of these $x$ values indicates that there is no strong configuration mixing. Such mixing would result in eigenstates made from chaotic superpositions of all configurations close in energy and smooth distribution of $x$ values. Instead, the distribution of $x$ consists of a small group of large values that dominate the average $\langle x\rangle$, and a broad distribution of smaller values. Thus, we cannot apply statistical theory developed for chaotic eigenstates [27, 28, 29, 30, 31].
@@
The results of our calculations for several possible transitions are presented in Table 1. Our findings show that the probability of nuclear excitation can, in principle, be enhanced 5 million times. However, the most probable cases correspond to $\beta(x_{m})$, presented in the last column of the table. In these cases, the enhancement may reach five orders of magnitude when the magnetic dipole hyperfine interaction is allowed (for $J_{n}=3/2$). This enhancement can be achieved through proper selection of laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
@@ IV Numerical calculation of the electronic bridge contribution to the decay of the nuclear excited state
Table 2: The values of $\beta$ for the EB contribution to the decay rate of the isomeric nuclear state. The rate depends on the value of the unknown total angular momentum $J_{n}$ of the state which is closest in energy to the energy of nuclear transition $\omega_{N}$ = 67393.34 cm-1. Possible values of $J_{n}$ for $\Delta=E_{n}-\omega_{N}=14.73$ cm-1 are 3/2, 5/2 and 7/2. If $J_{n}=7/2$, then the magnetic dipole hfi does not contribute, contribution of the the electric quadrupole hfi is small ($\beta=2$) and the leading contribution comes from the next closest state with energy $E_{n}$ = 67509.63 cm-1, $\Delta=E_{n}-\omega_{N}=116.29$ and $J_{n}=3/2$ (spectra from Ref.  [23]). $x\equiv\langle i||T_{k}||n\rangle^{2}\langle n||D||f\rangle^{2}$; $\langle x\rangle$ is the average value, $x_{m}$ is the median value.
$J_{n}$ $E_{n}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
[cm-1] [cm-1]
3/2 67378.61 14.73 3963 337
5/2 67378.61 14.73 5595 341
7/2 67378.61 14.73 2111Only electric quadrupole hfi contributes.
3/2 67509.63 116.29 63222Contribution from the next state $E_{n}$=67509.63  cm-1, in case of $J_{n}$=7/2 for $E_{n}$=67378.61 cm-1. 5
@@
The electronic bridge contribution to the decay of the nuclear excited state is described by the diagram in Fig. 2. This process requires summing over all possible final electronic states $f$. The sum over the intermediate states $n$ is dominated by a single term with energy $E_{n}=67378.61$ cm-1, which results in a small energy denominator in the hyperfine interaction induced EB amplitude: $\Delta=E_{n}-\omega_{N}=14.73$ cm-1.
Unfortunately, the electron angular momentum for this level is unknown, so we have considered all three possible cases: $J_{n}=3/2$, 5/2, and 7/2. When $J_{n}=7/2$, only the electric quadrupole hfi contributes. A larger magnetic dipole hfi contribution arises from a higher level with $J_{n}=3/2$, at $E_{n}=67509.63$ cm-1, with $\Delta=E_{n}-\omega_{N}=116.29$ cm-1.
We cannot assign definite wave functions to the states $n$ and the high-energy final states $f$. Therefore, we calculate the matrix element product $x\equiv\langle i||T_{k}||n\rangle^{2}\langle n||D||f\rangle^{2}$ for the lowest state $f$ and approximately 30 possible assignments of states $n$. The average value $\langle x\rangle$ and the median value $x_{m}$ are used in the calculation of the enhancement factors $\beta$ The contribution of the higher $f$ states is estimated assuming the $\omega^{3}$ dependence of the individual terms (see Eq. (2) ). Approximately 50 experimental energy levels are used in the estimation. The saturation of the summation was clearly observed.
We present the enhancement factor $\beta$ for both the average value $\langle x\rangle$ and the median value $x_{m}$ in Table 2.

### v2

@@ I Introduction
The amplitude of the nuclear M1 transition is suppressed by five orders of magnitude compared to typical atomic E1 transitions. However, it is possible to enhance this transition using the electronic bridge (EB) process, where electronic transitions induce nuclear transitions via magnetic interaction between electron and nucleus. The EB effect also increases the probability of decay of the nuclear excited state [14]. The high efficiency of excitation of the low-lying isomeric level in the 229Th nucleus by photons through the EB was predicted about 30 years ago in the works of Refs. [15, 16, 17]. More calculations have been performed for Th II, Th III, and Th IV ions, see e.g.  [18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Our previous EB calculations for the decay [19] and excitation [20] of the excited nuclear state in the Th II ion could not provide accurate results due to a large uncertainty in the measured value of the nuclear excitation energy and the absence of measurements of the positions of Th II electron energy levels near the nuclear excitation energy. According to theoretical expectations [28], the spacing between electron energy levels in this region is very small, making resonance with the nuclear excitation possible.
@@
II Electronic bridge for nuclear excitation in Th II
We follow the method used in Refs. [20, 25] for the Th II anf Th III ions. The decay of any atomic state with energy $E_{n}>\omega_{N}$ may include nuclear excitation. Since the nuclear excitation energy $\omega_{N}=67393$ cm-1 lies outside the optical region, we consider a two-step excitation of the electronic states(see also [20]). In the first step, the atom is excited from the ground state (GS) to one of the intermediate states $t$ used in [20]. In the second step, the ion is further excited by a second laser with frequency which is chosen to satisfy the energy conservation condition for simultaneous nuclear excitation and excitation of the final electronic state, $\omega_{1}+\omega_{2}=\omega_{N}+E_{s}$ ($\omega_{1}=E_{t}$). Nuclear excitation is due to the hyperfine interaction between atomic electrons and the nucleus. Corresponding diagram is presented in Fig. 1. Its calculation involves a summation over all intermediate states $n$. However, in this analysis, we focus on the dominant contributions, which come from specific intermediate states $n$ and final states $s$ providing a very small energy denominator in Eq. (3). Note that the final state $s$ is not the ground state but a low-lying excited state satisfying selection rules for the E1 transition induced by magnetic interaction between electron and nucleus, from the state $t$ (see diagram in Fig. 1). Suggestion that EB excitation process with final electron in an excited state has higher probability was made in Ref.  [29].
We do not go to highly excited states $s$ to keep laser frequencies within optical region. For each final electron state $s$, we look for an intermediate state $n$ (using energy level data from Ref. [31]) which corresponds to the smallest energy denominator $\Delta$ in (3), $\Delta=E_{n}-E_{s}-\omega_{N}$, and use this state for estimation of the EB amplitude.
@@ III Numerical calculations of the electronic bridge contribution to the nuclear excitation
Table 1: Versions of the EB process with resonant nuclear excitation. The notations $t$, $n$, and $s$ refer to the diagram in Fig. 1. It is assumed that $E_{t}$ = 36390.53 cm-1, $\omega_{2}=\omega_{N}+E_{s}-E_{t}$ is the frequency of second laser excitation. Final states $s$ are listed with their experimental energies ($E_{s}$) and calculated lifetimes ($\tau_{s}$). The energy denominator in Eq. (3) is given by $\Delta=E_{n}-E_{s}-\omega_{N}$. The third column lists all possible values of $J_{n}$ from the experimental data in Ref. [31], the last two columns present the calculated values of the EB enhancement factors $\beta$ for specific $J_{n}$. Here, $x\equiv\langle s||T_{k}||n\rangle^{2}\langle n||D||t\rangle^{2}$, and $\langle x\rangle$ represents the average value of $x$ over theoretical states that are candidates to represent the actual states $t$ and $n$. No averaging is performed over $s$ since it is a low-lying state that is easily identified in the calculations. The median value of $x$ is denoted by $x_{m}$. Numbers in square brackets indicate powers of ten.
$N$ $E_{n}$ [^1] Ref. [31], Supplemental Material. $J_{n}$ State $s$ $E_{s}$222 NIST database, [32]. $J_{s}$ $\tau_{s}$ $\omega_{2}$ $\Delta$ $\beta(\langle x\rangle)$ $\beta(x_{m})$
@@
1 73637.54 3/2 or 5/2 $6d^{2}7s$ 4P1/2 6244.29 1/2 7.8 s 37247.10 -0.09 5.5[6]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole interaction between electron and nucleus. 1.6[5]333 if $J_{n}$=3/2 for $E_{n}$=73637.54 cm-1, the dominant contribution comes from the magnetic dipole interaction between electron and nucleus.
2 73637.54 3/2 or 5/2 $6d^{2}7s$ 4P1/2 6244.29 1/2 37247.10 -0.09 1.2[4]444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole interaction between electron and nucleus contributes. 660444 if $J_{n}$=5/2 for $E_{n}$=73637.54 cm-1, only electric quadrupole interaction between electron and nucleus contributes.
3 74396.40 3/2 or 5/2 $6d^{3}$ 3/2 7001.42 3/2 110 ms 38004.23 1.64 2.5[4] 800
4 76788.94 1/2 or 3/2 $6d^{3}$ 4F5/2 9400.96 5/2 1.5 ms 40403.77 -5.36 5.6[4]555 if $J_{n}$=3/2. 700555 if $J_{n}$=3/2.
5 76895.40 3/2 $6d^{3}$ 4F5/2 9400.96 5/2 40403.77 101.1 3.0[3] [^6] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1. 4666Leading contribution if $J_{n}$=1/2 for $E_{n}$=76788.94 cm-1.
6 76445.96 1/2 or 3/2 $6d^{2}7s$ 4P5/2 9061.10 5/2 0.7 ms 40063.91 -8.48 7.1[3]555 if $J_{n}$=3/2. 24555 if $J_{n}$=3/2.
7 76508.84 3/2 - 7/2 $6d^{2}7s$ 4P5/2 9061.10 5/2 40063.91 54.4 1.1[3] [^7] Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1. 4777Leading contribution if $J_{n}$=1/2 for $E_{n}$=76445.96 cm-1.
8 68921.30 3/2 or 5/2 $6d^{2}7s$ 4F5/2 1521.90 5/2 12 s 32524.71 6.06 2.6[3] 28
9 75840.96 3/2 or 5/2 $6d^{3}$ 3/2 8460.35 5/2 2.6 ms 39463.16 -12.73 5.9[3] 94
@@
In the first step, we identified the final electronic state $s$ and the intermediate electronic level $n$ that provide very small energy denominators $\Delta=E_{n}-E_{s}-\omega_{N}$ in Eq. (3). The smallest value of $\Delta$ is $-0.09$ cm-1. We also considered five additional cases with larger $\Delta$, where enhanced electronic bridge contributions may still be expected.
@@
For each final state $s$, we calculated 217 possible values of $x$, corresponding to different assignments of the experimental levels $t$ and $n$. The distribution of these $x$ values indicates that there is no strong configuration mixing. Such mixing would result in eigenstates made from chaotic superpositions of all configurations close in energy and smooth distribution of $x$ values. Instead, the distribution of $x$ consists of a small group of large values that dominate the average $\langle x\rangle$, and a broad distribution of smaller values. A similar conclusion was reached in Ref. [35]. Thus, we cannot apply statistical theory developed for chaotic eigenstates [36, 37, 38, 39, 40].
@@
The results of our calculations for several possible transitions are presented in Table 1. Our findings show that the probability of nuclear excitation can, in principle, be enhanced 5 million times. However, the most probable cases correspond to $\beta(x_{m})$, presented in the last column of the table. In these cases, the EB enhancement factor $\beta$ may reach five orders of magnitude when the magnetic dipole hyperfine interaction is allowed (for $J_{n}=3/2$). This enhancement can be achieved through proper selection of laser frequencies in a two-step atomic excitation process, followed by nuclear excitation.
However, as it was noted in Ref. [27], a very large value of EB enhancement factor $\beta$ does not necessarily lead to the $\beta$ times enhancement of the excitation rate. The point is that the photon capture resonance cross section contains total width in denominator, which may be dominated by the radiative width $\Gamma_{\gamma}$, if it is very strongly enhanced by the EB process ($\Gamma_{\gamma}=(1+\beta)\Gamma_{N}$, where $\Gamma_{N}\approx 5\times 10^{-4}$ Hz is the the decay width of the nuclear isomeric state in bare nucleus - see e.g [42]). This effect can be qualitatively described by the factor, which appears in the resonance photon capture cross section
$\sigma\sim\lambda^{2}\frac{\Gamma_{\gamma}\Gamma}{(\omega-\omega_{0})^{2}+\Gamma^{2}/4},$ (7)
where $\lambda$ is the photon wavelength, $\Gamma=\Gamma_{\gamma}+\Gamma_{Ni}+\Gamma_{s}+\Delta_{l}$, $\Gamma_{Ni}$ is width of the nuclear transition in 229mTh+ , $\Gamma_{s}$ is the width of the final electron state and $\Delta_{l}$ is the width of the energy distribution of the photons in the pumping laser beam (or width of the frequency comb teeth). There may be other line broadening mechanisms.
In the resonance $\sigma\propto\beta\Gamma_{N}/(\beta\Gamma_{N}+\Gamma_{Ni}+\Gamma_{s}+\Delta_{l})$. We see that in the limit $\beta\to\infty$ the enhancement in the center of the resonance disappears.
However, in a realistic situation there is $\beta$ times enhancement of the cross section. Consider different contributions to the total width $\Gamma$. Specifically, for the final electron states 4P1/2 with the energy 6244.29 cm-1 and 5F5/2 with the energy 1521.90 cm-1 , the radiative width $\Gamma_{s}$ is very small, 0.1 Hz, since it is produced by low frequency M1 and E2 transitions. For other final states in the Table 1, E1 transitions are possible, and the width is from 10 to 1000 Hz. Note that all lifetimes are strongly affected by the configuration mixing.
There is an experimental indication that the lifetime of the nuclear isomer in 229mTh+ with electrons in their ground state is smaller than 10 ms [41] (see discussion in the next section), i.e. in 229mTh+ the natural decay width $\Gamma_{Ni}>$ 100 Hz. When electrons are in exited state, $\Gamma_{Ni}$ would hardly be smaller. If this is the case, there is an enhancement of the cross section at least up to $\beta=2\times 10^{5}$.
Moreover, during the frequency scanning procedure we are not exactly in the resonance, i.e. $(\omega-\omega_{0})^{2}>\Gamma^{2}/4$. The width of the frequency comb teeth $\Delta_{l}$ may significantly exceed natural width $\Gamma$. Finally, the scanning time is reduced inversely proportional to the width, so effectively one should consider the product of the cross section times the total width, so $\beta$ in the denominator cancels out. In this case the EB enhancement factor $\beta$ is important for any large $\beta$.
@@ IV Numerical calculation of the electronic bridge contribution to the decay of the nuclear excited state
Table 2: Calculations of the coefficient $R$ (8) at different assumptions for experimental level at $E_{n}$=67378 cm-1 ($J_{n}$=3/2,5/2 or 7/2). $N$ is serial number of the state in the calculations for given $J_{n}$ and parity; $\Delta E_{n}$ is the shift in the energy of state $n$ to bring it to experimental energy of 67378 cm-1. Numbers in square brackets stand for powers of ten.
$N$ $E_{n}$ cm-1 $\Delta E_{n}$ Partial $J_{f}$ contributions Sum
$J_{n}$=3/2; $J_{f}:$ 1/2 3/2 5/2
68 67601 222 8.524[1] 1.414[2] 8.824[1] 3.149[2]
69 67914 536 1.908[3] 3.929[3] 1.695[3] 7.532[3]
70 68174 796 4.276[2] 5.208[2] 2.036[2] 1.152[3]
71 68568 1189 6.270[3] 3.274[3] 1.933[3] 1.148[4]
72 68866 1488 2.554[2] 2.161[2] 2.261[2] 6.976[2]
73 69131 1752 6.223[2] 6.795[2] 1.279[3] 2.580[3]
74 69391 2013 1.762[3] 1.398[3] 2.280[3] 5.440[3]
75 69500 2121 1.548[3] 4.602[3] 4.080[3] 1.023[4]
76 69719 2340 8.695[3] 6.655[3] 1.121[4] 2.655[4]
77 69874 2496 5.716[2] 5.475[2] 6.200[2] 1.739[3]
78 70221 2842 5.813[2] 2.826[3] 1.475[3] 4.882[3]
79 70583 3204 2.090[3] 6.944[3] 8.203[3] 1.724[4]
80 71018 3639 1.665[3] 1.943[3] 3.999[3] 7.607[3]
81 71104 3725 2.949[2] 3.232[2] 5.333[2] 1.151[3]
82 71270 3892 3.868[3] 4.111[3] 9.680[3] 1.766[4]
83 71528 4149 1.113[2] 6.798[1] 1.302[2] 3.095[2]
84 71816 4437 1.108[3] 7.471[2] 4.730[3] 6.586[3]
85 72150 4771 2.141[3] 4.601[3] 1.353[4] 2.028[4]
$J_{n}$=5/2: $J_{f}:$ 3/2 5/2 7/2
78 67493 115 1.624[4] 2.049[4] 4.874[3] 4.160[4]
79 67915 536 9.062[2] 4.504[2] 2.893[2] 1.646[3]
80 68350 971 2.976[4] 1.724[4] 2.396[4] 7.096[4]
81 68531 1153 3.824[4] 5.077[4] 4.200[4] 1.310[5]
82 69006 1628 5.328[2] 1.447[3] 7.728[2] 2.752[3]
83 69196 1818 1.458[4] 3.942[4] 3.654[4] 9.054[4]
84 69309 1930 6.857[3] 5.275[3] 2.767[3] 1.490[4]
85 69352 1973 9.055[2] 1.242[3] 6.894[2] 2.837[3]
86 69640 2261 8.408[3] 2.196[4] 1.227[4] 4.264[4]
87 69817 2439 6.820[3] 2.231[4] 1.243[4] 4.155[4]
88 69918 2539 3.727[2] 8.864[2] 6.485[2] 1.908[3]
89 69974 2596 6.553[2] 1.908[3] 2.102[3] 4.665[3]
90 70222 2844 1.637[3] 3.367[3] 2.770[3] 7.774[3]
91 70475 3096 5.853[3] 6.302[3] 3.767[3] 1.592[4]
92 70510 3131 9.887[3] 9.563[3] 5.978[3] 2.543[4]
93 70823 3444 9.029[1] 1.072[2] 6.852[1] 2.660[2]
94 70924 3546 8.687[2] 9.090[2] 5.178[2] 2.296[3]
95 71035 3656 2.439[2] 2.602[2] 1.229[2] 6.270[2]
@@
The electronic bridge contribution to the decay of the nuclear excited state is described by the diagram in Fig. 2. This process requires summing over all possible final electronic states $f$. The sum over the intermediate states $n$ is dominated by a single term with energy $E_{n}=67378.61$ cm-1 [31], which results in a small energy denominator in EB amplitude induced by magnetic dipole or electric quadrupole interaction with the nucleus: $\Delta=E_{n}-\omega_{N}=14.73$ cm-1. Very little is known about this state, only energy and a range of possible values of the total angular momentum ($J_{n}$=3/2, 5/2 or 7/2). Therefore, we consider a number of possible candidates (see below).
As before, it is convenient to present the results in terms of dimensionless ratios of EB transition rates to the nuclear transition rate. Note, that summation over final states $f$ leads to summation of terms similar to $\beta_{M1}$ (5) and $\beta_{E2}$ (6) with different values of $\omega$: $\omega\equiv\omega_{f}=\omega_{N}-E_{f}-E_{i}$. Therefore, we introduce new variable $R$
$R=\sum_{f}\beta(\omega_{f}).$ (8)
In the paper  [41] they study internal conversion signal which should appear after collision of the 229mTh isomer ions with a solid surface. They see clear signals from Th III and Th IV ions but do not observe any signal from Th II ions. The authors of Ref.  [41] suggested a possible explanation: lifetime ($\tau$) of the 229mTh isomer is very small, $\tau<10~{}{\rm ms}$, so Th II ions do not reach the surface.
The half-life of the isomeric nuclear state in Th IV is practically not affected by the electronic bridge mechanism [18]. This half-life in Th IV was measured to be $T_{1/2}=1400^{+600}_{-300}$ [42], corresponding to the lifetime $\tau=2020^{+866}_{-433}$. This lifetime is in a reasonable agreement with the values obtained in the solid state experiments [11, 12, 13, 43, 47, 44, 45, 46, 48] after introducing the solid state corrections. To obtain $\tau<10~{}{\rm ms}$, one needs very large $R>2\times 10^{5}$. This small value of the lifetime constitutes the so called thorium puzzle. Possible solutions of this puzzle (very large value of EB coefficient, $R=56500$, collisional broadening) were discussed in Refs.  [35, 29, 30].
We study this problem in present work using new experimental data on Th+ energy levels [31]. Strong enhancement of the decay rate due to above mentioned small energy denominator ($\Delta=14.73$ cm-1) may potentially explain the puzzle. We perform the calculations summing over large number of intermediate states $n$ around $E_{n}\sim\omega_{N}$ (about 20 states for each $J_{n}$) and all final states with $E_{f}<\omega_{N}$ (about 70 states for each $J_{f}$).
Comparison with experiment for excitation energies between 0 and 60000 cm-1 shows that when approaching the upper limit 60000 cm-1, calculated energies tend to be about 3000 cm-1 higher than experiment. Since we need states with energies $E_{n}\sim\omega_{N}$ = 67393.34 cm-1, it does make sense to shift calculated energies down to make more realistic energy denominators.
[…]
The largest values of $R$ in the Table, $R=1.3\times 10^{5}$, corresponds to the assumption that the experimental energy $E_{n}$=67378.61 cm-1 match up with the calculated state number 81 with the calculated energy 68531 cm-1 and total angular momentum $J_{n}=5/2$. This maximal calculated value of $R$, $R=1.3\times 10^{5}$, is close to the experimental lower bound, $R>2\times 10^{5}$. Therefore, we cannot entirely exclude this scenario as a potential solution to the thorium puzzle: an unexpectedly large enhancement of the nuclear decay rate via the electronic bridge mechanism [35, 29, 30].
We have also performed calculations for many levels located below nuclear frequency $\omega_{N}$ = 67393.34 cm-1, but they do not produce any qualitative changes to the picture.